OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

when two lines are parallel their slopes are equal thus the slope from the given equation is 7. here you got slope =7 and a point (1,-3) use point slope form to get the equation point slope form is y-y1 =m(x-x1) here m=7 , y1=-3 , x1= 1 solve it now

OpenStudy (anonymous):

is it a or c?

OpenStudy (anonymous):

what do you think??

OpenStudy (anonymous):

Ehhh if I knew I'd know the answer -_-

OpenStudy (anonymous):

A